Find the partial sum Sn of the geometric sequence that satisfies....1+4+16+64...+262144
Answer by stanbon(75887)

Find the partial sum Sn of the geometric sequence that satisfies....1+4+16+64...+262144
---
Note: d = 3
Need to find "n"
a(n) = a(1)+(n-1)d
-------------------------
262144 = 1 + (n-1)3
262143 = 3n-3
262146 = 3n
n = 87382
-------------------------
So there are 87382 terms in the sequence.
-------------------------
S(n) = (n/2)(a(1)+a(n))
----
S(87382) = (87382/2)(1+262144)
--
I'll let you do the arithmetic.
Cheers,
Stan H.